OpenStudy (anonymous):

Use Mathematical Induction to prove the statement is true for all positive integers n, of show why it is false. 1^2+4^2+7^2+...+(3n-2)^2=n(6n^2-3n-1)/2

OpenStudy (anonymous):

did you check to see if it was true first ?

OpenStudy (anonymous):

I dont know how to do that...

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

Yeah so you see my dilemma

OpenStudy (anonymous):

\[\sum_{k=1}^n(3k-2)^2\] is what you have right?

OpenStudy (anonymous):

What?? how did you get that?

OpenStudy (anonymous):

i wrote your sum in sigma notation is all

OpenStudy (anonymous):

oh okay haha so how do i do this problem?

OpenStudy (anonymous):

to check that it is indeed true, note that \[(3k-2)^2=3k^2-12k+4\] so \[\sum_{k=1}^n(3k-2)^2=3\sum_{k=1}^n k^2 -12\sum_{k=1}^nk +4\sum_{k=1}^n 1\]

OpenStudy (anonymous):

you have formulas for each of these three, so it it is algebra from here on in that is how your prof made up the problem

OpenStudy (anonymous):

okay so using algebra how would i prove this? Do i just plug in the numbers to k?

OpenStudy (anonymous):

you use the following \[\sum_{k=1}^nk^2=\frac{n(n+1)(2n+1)}{6}\] \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] \[\sum_{k=1}^n1=n\]

OpenStudy (anonymous):

\[\sum_{k=1}^n(3k-2)^2=3\sum_{k=1}^n k^2 -12\sum_{k=1}^nk +4\sum_{k=1}^n 1\] \[=3\left(\frac{n(n+1)(2n+1)}{6}\right)-12\left(\frac{n(n+1)}{2}\right)+4n\]

OpenStudy (anonymous):

or \[\frac{n(n+1)(2n+1)}{2}-6n(n+1)+4n\] i will let you finish the algebra for your self then we can prove it is true by induction

OpenStudy (anonymous):

I am so confused.

OpenStudy (anonymous):

yes i sensed that that was true. lets back up a bit are you familiar with this sigma notation?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

oh okay then lets forget it

OpenStudy (anonymous):

\[1+2+3+4+...+n=\frac{n(n+1)}{2}\] how about that formula? have you seen it?

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

not before now no

OpenStudy (anonymous):

ok then lets forget that also, although that is a very good formula to know it tells you how to add up the first consecutive \(n\) whole numbers, like \[1+2+3+...+99+100=\frac{100\times 101}{2}=5050\]

OpenStudy (anonymous):

okay i get that formula

OpenStudy (anonymous):

however, if you do not know these formulas, then we cannot do step one, which is to show that that \[1^2+4^2+7^2+...+(3n-2)^2=\frac{n(6n^2-3n-1)}{2} \] is correct

OpenStudy (anonymous):

so lets just assume that it is correct, and proceed to prove it by induction ready?

OpenStudy (anonymous):

Is there any way you could just give me the answer and work backwards to show how you found it?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

first of all you have to check that it is true if \(n=1\)

OpenStudy (anonymous):

This equation does not equal eachother

OpenStudy (anonymous):

if n=1

OpenStudy (anonymous):

so you have to replace \(n\) by 1 on the right side of the equal sign that is, you have to show that \[1=\frac{1\times (6\times 1^2-3\times 1-1)}{2}\] is it clear why i wrote that?

OpenStudy (anonymous):

yes i get that part

OpenStudy (anonymous):

ok and if it does not work, you are done, it is not true but it does work, do the arithmetic carefully and you will see it is right

OpenStudy (anonymous):

well i got 1 on the left side and 2 on the right side

OpenStudy (anonymous):

did you remember to divide by 2?

OpenStudy (anonymous):

No! so they are equal! what do i do now?

OpenStudy (anonymous):

\[1=\frac{1\times (6\times 1^2-3\times 1-1)}{2}=\frac{6-3-1}{2}=\frac{2}{2}=1\]

OpenStudy (anonymous):

Okay i get that part. Its pretty easy

OpenStudy (anonymous):

ok now comes the tricky part. we get to "assume it is true if \(n=k\)" so lets replace \(n\) by \(k\) and write what we know, namely that \[1^2+4^2+7^2+...+(3k-2)^2=\frac{k(6k^2-3k-1)}{2}\]

OpenStudy (anonymous):

i hope you see i did almost nothing, just replaced \(n\) by \(k\) everywhere i saw it. now we have to prove that it is true if \(n=k+1\) the first thing we need is to know what to write if \(n=k+1\) have you done a proof by induction before?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

then this is a really bad one to start with because the algebra is going to suck have you SEEN a proof by induction before?

OpenStudy (anonymous):

no haha and just my luck i guess

OpenStudy (anonymous):

hmm, then i think this is not going to make the slightest bit of sense to you do you have to hand this in? is it for a class?

OpenStudy (anonymous):

yeah its a worksheet that i have to scan into my computer and send to my teacher. it is an online math class.

OpenStudy (anonymous):

Is there any way that you can just give me the answer so that i can finish the worksheet?